A two-slit interference experiment is set up, and the fringes are displayed on a screen.

Then the whole apparatus is immersed in the nearest swimming pool. How does the fringe pattern change?



1
Expert's answer
2022-02-20T15:49:16-0500

Position of "m^{th}" bright fringe in air medium,


"y_m=R\frac{m\lambda_\omicron}{d}"..... (1)


Position of "m^{th}" bright fringe in water medium,


"{y_m}'=R\frac{m\lambda_W}{d}"..... (2)


in equation (1) and (2)

R is the distance from slits to screen

m is the order of the fringe

"\lambda_W" is the wavelength of the light in water medium

"{\lambda_\omicron}" is the wavelength of the light in air medium

d is the distance between slits


Divide equation (II) with equation (I),

"\frac{{y_m}'}{{y_m}}=\frac{R\frac{m\lambda_W}{d}}{R\frac{m\lambda_\omicron}{d}}"

"=\frac{\lambda_W}{\lambda_\omicron}"


The wavelength of the light in water medium is given by,

"\lambda_W=\frac{\lambda_\omicron}{n}"

where refractive index of the water is n,


Substitute "\frac{\lambda_\omicron}{n}" for "\lambda_W" to find "{y_m}'" ,

"\frac{{y_m}'}{{y_m}}=\frac{\frac{\lambda_\omicron}{n}}{\lambda_\omicron}"

"=\frac{1}{n}"

Substitute 1.33 for n to find "{y_m}'" ,

"\frac{{y_m}'}{{y_m}}=\frac{1}{1.33}"

"=\frac{3}{4}"

"{y_m}'=\frac{3{y_m}}{4}"


Therefore, the new fringe pattern is "\frac{3}{4}" times the old fringe pattern.